If 10.4 mols of hydrogen, H2, and 9.1 mols of oxygen, O2, are placed together in a container and allowed to react according to the equation above, which one, if any, is in excess and by how much? also, how many moles of water, H2O, would be produced by the reaction?


1
Expert's answer
2021-03-04T06:41:18-0500

Solution:

2H2 + O2 = 2H2O

10,4mol   9,1mol

xmol(O2)=10,4/2=5,2

9,1-5,2=3,9
surplus(O2)= 3,9mol

Excess (O2)=3,9mol

xmol(H2O)=10,4*2/2=10,4

Answer:

surplus(O2)= 3,9mol

H2O=10,4 mol